OpenStudy (anonymous):

What are the coordinates of the center of the ellipse shown below? (x+2)^2 (y-4)^2 ------- + -------- = 1 9 36 (2,-4) (-2,4) (9,36) (3,6)

OpenStudy (whpalmer4):

when you write the ellipse in that form, (x-h)^2/a^2 + (y-k)^2/b^2 = 1 the center is at (h,k), and the axes are of length a and b if a^2 > b^2, then it is horizontal, otherwise vertical if a^2 < b^2. if a^2=b^2 it is the special case of an ellipse known as a circle :-)

OpenStudy (whpalmer4):

sorry, the semi axes...

OpenStudy (anonymous):

so what do I do next?

OpenStudy (whpalmer4):

well, all you have to do is compare your equation with the form I gave you, and read out the coordinates of the center

OpenStudy (whpalmer4):

you have (x+2)^2/9 + (y-4)^2/36 = 1 I have (x-h)^2/a^2 + (y-k)^2/b^2 = 1 what is the value of h? what is the value of k? for extra credit, what are the values of a and b? when you know h and k, you know the center is at (h, k)

OpenStudy (anonymous):

-2,4?

OpenStudy (anonymous):

a is 3 and b is 6

OpenStudy (whpalmer4):

if h = -2, and k = 4, the center is at (-2, 4)

OpenStudy (whpalmer4):

that was easy, right? :-)

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

Sometimes it is more complicated, and you have to do some algebra to get it into that form. For example, y^2 - 8y + 4x^2 + 16x - 4 = 0 is actually an equation for that same ellipse, but it sure isn't obvious, is it?

OpenStudy (whpalmer4):

but by shuffling things around and completing the square on both x and y, you could get it into that same form we had and read out the center and semi-major/minor axes